The standard form equation for the ellipse with major axis endpoints (+ or - 5, 0) and minor axis length 4 is? Answer by lwsshak3(11628) 
 The standard form equation for the ellipse with major axis endpoints (+ or - 5, 0) and minor axis length 4 is.
**
Standard form of an equation for an ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
..
For given equation:
center:(0,0)
given length of major axis=10=2a
a=5
a^2=25
..
given length of minor axis=4=2b
b=2
b^2=4
..
Equation:
x^2/25+y^2/4=1